TRUE OR FALSE QUESTION please explain the reason briefly (3 sentences)

3) Max’s utility function is U(x,y) = (x+2)(y+1). If his consumption of both x and yare doubled, then his marginal rate of substitution between x and y remains constant.
Answer #1

False, we can derive marginal rate of subsitution in below and we can say that, if values are doubled for X and Y, MRS will change and it is not a constant.

False, Given utility fun tion; u(x, y) = (x + 2) (y11) U (x, y) = xy + x + 2y + 2 we know that Morgenal rate of substitution;